the perimeter of a rectangle is 44 inches.
what is the are of the rectangle?

Answer by Alan3354(69443)

the perimeter of a rectangle is 44 inches.
what is the are of the rectangle?
-------------
P = 2L + 2W = 44
L + W = 22
Area = L*W
----------
Area = L*(22 - L)
L can be any number 0 < L < 22